# Knowledge-based Feedback Generation

### v1

@@ Abstract
Humans follow criteria when they execute tasks, and these criteria are directly used to assess the quality of task completion. Therefore, having models learn to use criteria to provide feedback can help humans or models to perform tasks better. However, existing research in this field tends to consider only a limited set of criteria or quality assessment aspects. To fill this gap, we propose a general framework that enables large language models (LLMs) to use comprehensive criteria for a task in delivering natural language feedback on task execution. In particular, we present a model-in-the-loop framework that semi-automatically derives criteria from collected guidelines for different writing tasks and constructs in-context demonstrations for each criterion. We choose three tasks from real-world scenarios to operationalize this idea: paper introduction writing, Python code writing, and Reddit post writing, and evaluate our feedback generation framework using different LLMs. The results reveal the fine-grained effects of incorporating criteria and demonstrations and provide valuable insights on how to teach LLMs to use criteria more effectively.
@@ 1 Introduction
    Experiment results suggests that providing criteria allows the model to generate feedback that contains more critiques and suggestions. In addition, providing demonstrations makes the resulting critiques and suggestions more helpful, but may also distract the model into generating feedback on the demonstration input.
@@
    We release 83 criteria and 332 in-context demonstrations that we have collected and curated for the three real-world writing tasks at https://github.com/yyy-Apple/LLMCrit.
@@ 3.3 Create Demonstrations
Create Demonstration Input To increase the diversity of demonstrations, we aim to collect input texts that randomly violate some criteria while complying with others in an aspect. For an aspect $a_{i}$ with criteria $c_{i}=\{c_{i}^{1},\cdots,c_{i}^{N}\}$, we first sample a satisfaction vector $s_{i}=\{s_{i}^{1},\cdots,s_{i}^{N}\}$ where $s_{i}^{j}=1$ indicates satisfaction of criterion $c_{i}^{j}$ and $s_{i}^{j}=0$ the violation of $c_{i}^{j}$. Then, we prompt an LLM with $a_{i}$, $c_{i}$, and $s_{i}$ to get a candidate input text $x_{i}$, followed by manual refinement to ensure compliance with $s_{i}$. We repeat this process $k$ times to collect $k$ demonstrations for $a_{i}$. One point worth noting is that our preliminary experiments suggest that by relying solely on $a_{i}$, $c_{i}$, and $s_{i}$ to create the input text, LLMs (even the most advanced ones) are likely to generate text that contains unnatural and simplistic errors. To address this problem, we use existing human-written texts $\tilde{x}_{i}^{1}\cdots\tilde{x}_{i}^{k}$ as a starting point, which are then modified by the LLM.
Create Demonstration Output To get the demonstration output $y_{i}^{j}$, we prompt an LLM with $x_{i}^{j}$, $a_{i}$ and $c_{i}$to get the initial feedback. Subsequently, human experts would refine this output, addressing minor issues such as factual errors and enhancing its structure and clarity.
@@ 4.5 Evaluation Methodology
The most straightforward way to judge the quality of the model-generated feedback is to measure whether it provides helpful critiques (throughout the paper, critiques refer to critiques that point out flaws) or suggestions for each criterion. Given the resource-intensive nature of human evaluation and recent findings indicating that LLMs can assess model outputs in a manner that aligns with human evaluations (Li et al., 2023b; Zheng et al., 2023b), we resort to model-based evaluation. We employ Claude2 for evaluation due to its free research accessibility and support for large context length. However, assessing helpfulness in one step still poses a great challenge to Claude2 according to our preliminary experiments due to various factors that could make a feedback text unhelpful. Therefore, we propose a layered evaluation strategy and decompose the “helpfulness” evaluation into four progressive perspectives.
@@
    Validity Whether the generated text is a valid feedback text. Failures include offering general writing advice, merely echoing the criteria, or continuing the provided input text.
@@
    Contextualization Whether the generated text is a feedback text specific to the current input. Failures include when the generated feedback text is for a demonstration input.
@@
    Constructiveness Whether the feedback text provides critiques or suggestions w.r.t. a certain criterion. Failures include summarizing the text or providing exclusively positive comments.
@@
    Helpfulness Whether the feedback text contains critiques or suggestions that are helpful for improving the text writing w.r.t. a certain criterion. Failures include critiques or suggestions that are too vague or contain factual errors.
We designed different prompt templates for evaluating each perspective (see Appendix Tab. 13~Tab. 24) and validated our model-based evaluation approach by running a small-scale human evaluation conducted by two experienced NLP researchers using the same prompts as annotation guidelines. For each perspective and task, we randomly chose 100 feedback samples from all models, totaling 300 samples per perspective for annotation. The results indicate that our model-based evaluation method is highly accurate, with average accuracies of 96.48% for validity (95.33% inter-annotator agreement), 94.23% for contextualization (92% inter-annotator agreement), 95.22% for constructiveness (91% inter-annotator agreement), and 88.53% for helpfulness (89.67% inter-annotator agreement). [^7] When calculating the accuracy, we only include the samples where both annotators agreed on the label.
@@ 5 Experiments
In the case of the Crit, ICL, and CrICL strategies, one input text will receive multiple feedback texts, corresponding to different aspects.
@@ 6.1 Validity and Contextualization
Our study begins by examining LLMs’ capabilities in generating valid and contextual feedback texts for given examples. Specifically, we calculate the percentage of feedback texts that are valid and the percentage of valid feedback texts that are contextual. The detailed results are in Appendix Tab. 25. Our key findings are as follows: (i) Regardless of the teaching strategy, these LLMs almost always generate valid feedback texts, demonstrating their strong instruction-following and in-context learning capabilities. (ii) Including demonstrations can potentially distract models, leading them to write feedback for the input text of a demonstration, which can sometimes reduce the overall contextualization of the generated feedback texts. This risk is particularly high when the demonstrations are lengthy, such as in the introduction task.
@@ 6.2 Constructiveness and Helpfulness
For feedback texts that passed the validity and contextualization checks, we scrutinize their constructiveness by counting the percentage of criteria in §4 they touched upon by providing critiques or suggestions. When feedback texts contain critiques or suggestions for specific criteria, we assess their helpfulness by calculating the percentage of these criteria that can be improved using the corresponding critiques or suggestions. Results are in Tab. 1.
@@
(i) Across all models and tasks, including criteria generally significantly enhances the constructiveness of the generated feedback texts (on average by 25.8%, up to 44.6%), outperforming the base strategy. (ii) In most cases, adding demonstrations improves constructiveness of generated feedback texts compared to the base strategy (on average by 10.1%, up to 27.0%) . Only in a few cases, such as the introduction task with the Command and LongAlpaca models, does the addition of demonstrations lead to a noticeable decrease in constructiveness, likely due to struggles with extremely lengthy demonstrations. (iii) Interestingly, when both demonstrations and criteria are applied, the outcome is mostly comparable to adding demonstrations alone. This may be attributed to the extended length of demonstrations, which could potentially dilute the impact of the criteria.
@@
(i) The impact of adding criteria on helpfulness varies by model. For GPT4, adding criteria always helps to generate more helpful critiques and suggestions compared to the base strategy. For the other models, on the other hand, in all tasks, adding only criteria typically results in a marginally lower rate of helpful critiques and suggestions, as compared to the base strategy. (ii) In most cases, adding demonstrations could improve the helpfulness compared to the base strategy (on average by 2.7%, up to 13.7%). Exceptions occur when the base strategy already produces highly helpful critiques and suggestions (over 90% helpfulness). (iii) In most cases, combining both demonstrations and criteria yields inferior results compared to using demonstrations alone. This may be attributed to the criteria’s lack of specific guidance on crafting helpful critiques and its potential interference with the efficacy of demonstrations when added to the context.
@@ 6.3 Overall Performance
We also present the overall performance of each model using different teaching strategies. In particular, we count the total number of criteria the generated feedback texts have addressed by providing helpful critiques or suggestions for the 100 test samples for each task. The results are shown in Tab. 2. We find that (i) Including criteria consistently enhances overall performance across all models and tasks. (ii) Incorporating demonstrations generally improves overall performance, except when overly long demonstrations adversely affect the contextualization of feedback texts, as seen in the introduction task for most models. (iii) Combining both criteria and demonstrations typically does not outperform adding criteria alone, mostly because models generate fewer critiques and suggestions when demonstrations are provided compared to criteria only (see Appendix Tab. 26). This may be due to the fact that the feedback texts in demonstrations can include positive comments, leading to less critical model-generated feedback. Therefore, in practice, providing only criteria to LLMs is a reasonably effective strategy for feedback generation.
@@
Table 2: Overall performance of each model using different teaching strategies in terms of the number of criteria the generated feedback texts addressed through providing helpful critiques or suggestions.
@@ 7.1 Fine-grained Analysis of Criteria and Demonstration Effects
We selected the code task as our primary testbed, owing to its extensive number of criteria. For each criterion, we analyzed the proportion of generated critiques out of the total potential critiques that the model could produce for 100 test samples. By “potential critiques”, we mean that we assume all the feedback texts that passed the validity and contextualization checks contain critiques for all criteria in the aspect they were generated for. The percentage of generated critiques for each criterion is depicted on the left in Tab. 3. Furthermore, we examined the proportion of helpful critiques within the set of generated critiques for each criterion, with these percentages presented on the right in Tab. 3.
We observe that: (i) adding criteria enhances the model’s ability to generate critiques across various criteria. This enhancement is particularly effective for criteria that the model is already familiar with, as indicated by a higher baseline percentage of critiques. (ii) Adding demonstrations typically results in a more uniform distribution of helpful critiques across all criteria, indicating an overall improvement in critique quality. This approach contrasts with the base strategy, which tends to overlook certain criteria.
@@
To examine the impact of providing criteria at the aspect level versus the individual criterion level, we chose the Reddit post writing task, where each aspect has an average of 4.2 criteria. We only compared the feedback texts generated by the strategy of offering criteria only. The results are in Tab. 4. We see that providing criteria at the aspect level almost always positively affects the validity and contextualization of the generated feedback texts. However, it negatively affects the constructiveness of the generated feedback, while the impact on helpfulness varies by model. In summary, providing criteria at the aspect level, rather than at the individual criterion level, does not significantly alter overall performance, but it does offer greater efficiency during inference time.
@@ 8 Conclusion
In this paper, we present a general framework for teaching LLMs to use criteria for feedback generation. This framework takes the guidelines as a starting point, semi-automatically extracts criteria from them, and constructs in-context demonstrations. We then apply these criteria and in-context demonstrations to guide the feedback generation process of LLMs. In order to provide a more comprehensive assessment of the quality of generated feedback, we propose a layered evaluation methodology to measure the quality of generated feedback from different perspectives. Experiments conducted on three writing tasks and seven different LLMs provide insights into the most effective way of teaching LLMs to use criteria. We also examined the effectiveness of applying criteria at different levels of granularity and found that providing criteria at the aspect level is more effective than at the individual criterion level.
@@ 9 Limitations and Future Work
In our experiments, we focused on LLMs with strong instruction-following capabilities and large context windows. However, there are models that may not excel as much in instruction following and in-context learning as the models we selected. Exploring how to teach these models to use criteria would be an interesting avenue for future research. Additionally, our evaluation primarily relies on a model-based approach, which may introduce some inaccuracies, despite the meta-evaluation showing good correlation with human judgments. The development of improved model-based methods for measuring feedback quality is another important future research direction.
@@ A.5 Prompts For Layered Evaluation
A.6 Detailed Experiment Results
A.6.1 Results on Validity and Contextualization

### v2

@@ Abstract
Humans follow criteria when they execute tasks, and these criteria are directly used to assess the quality of task completion. Therefore, having models learn to use criteria to provide feedback can help humans or models to perform tasks better. However, current research in this area tends to consider only a limited number of criteria, or only a limited number of quality assessment aspects. To fill this gap, we propose a general framework that enables large language models (LLMs) to use comprehensive criteria for a task in delivering natural language feedback on task execution. In particular, we present a model-in-the-loop framework that semi-automatically derives criteria from collected guidelines for different writing tasks and constructs in-context demonstrations for each criterion. We choose three tasks from real-world scenarios to operationalize this idea: paper introduction writing, Python code writing, and Reddit post writing, and evaluate our feedback generation framework using different LLMs. The results reveal the fine-grained effects of adding criteria and demonstrations and provide valuable guidance on how to teach LLMs to use criteria more effectively.
@@ 1 Introduction
    Experiment results suggests that providing criteria allows the model to generate feedback that contains more critiques and suggestions. In addition, providing demonstrations makes the resulting critiques and suggestions more helpful, but may also distract the model from generating feedback on the demonstration input.
@@
    We release 83 criteria and 332 in-context demonstrations that we have collected and curated for the three real-world writing tasks to the community for future research. All resources can be found at https://github.com/yyy-Apple/LLMCrit.
@@ 3.3 Create Demonstrations
Create Demonstration Input To increase the diversity of demonstrations, we aim to collect input texts that randomly violate some criteria while complying with others in an aspect. For an aspect $a_{i}$ with criteria $c_{i}=\{c_{i}^{1},\cdots,c_{i}^{N}\}$, we first sample a satisfaction vector $s_{i}=\{s_{i}^{1},\cdots,s_{i}^{N}\}$ where $s_{i}^{j}=1$ indicates satisfaction of criterion $c_{i}^{j}$ and $s_{i}^{j}=0$ the violation of $c_{i}^{j}$. Then, we prompt an LLM with $a_{i}$, $c_{i}$, and $s_{i}$ to get a candidate input text $x_{i}$, followed by manual refinement to ensure compliance with $s_{i}$. We repeat this process $k$ times to collect $k$ demonstrations for $a_{i}$. One point worth noting is that our preliminary experiments suggest that by relying solely on $a_{i}$, $c_{i}$, and $s_{i}$ to create the input text, LLMs (even the most advanced ones) are likely to generate text that contains unnatural and simplistic errors. To address this problem, we use existing human-written texts $\tilde{x}_{i}^{1}\cdots\tilde{x}_{i}^{k}$ as a starting point. These texts are then modified by the LLM to introduce specific criteria violations.
Create Demonstration Output To get the demonstration output $y_{i}^{j}$, we prompt an LLM with $x_{i}^{j}$, $a_{i}$ $c_{i}$, and $s_{i}^{j}$ to get the initial feedback. Subsequently, human experts would refine this output, addressing minor issues such as factual errors and enhancing its structure and clarity.
@@ 4.5 Evaluation Methodology
The most straightforward way to judge the quality of the model-generated feedback is to measure whether it provides helpful (i.e., actionable and free of factual errors) critiques. [^7] Throughout the paper, critiques refer to violations of the criteria, or suggestions how to improve the writing with respect to those criteria. Given the resource-intensive nature of human evaluation and recent findings indicating that LLMs can assess model outputs in a manner that aligns with human evaluations (Li et al., 2023b; Zheng et al., 2023b), we resort to model-based evaluation. We employ Claude2 for evaluation due to its free research accessibility and support for large context length. However, our initial findings indicate that Claude2 struggles with assessing feedback helpfulness in a single step, as its judgments often do not correlate well with human evaluations. Upon detailed examination of the feedback generated by LLMs, we identified several key issues that contribute to the unhelpfulness of the feedback. These issues can be categorized into four primary types of errors:
@@
    Validity Error: The generated text does not qualify as feedback, but rather offers general writing advice, extends the given criterion, or simply repeats it.
@@
    Context Error: The feedback pertains to an in-context example rather than the specific example at hand.
@@
    Insufficient Critique: The feedback lacks depth and specificity. It may merely summarize the original text, provide broad and general comments without detailed critiques, or offer only positive remarks without any constructive suggestions.
@@
    Unhelpful Critique: The feedback includes critiques or suggestions that are either too vague to be actionable or are factually incorrect
In light of these challenges, we propose a refined evaluation methodology. We suggest implementing a layered evaluation strategy that breaks down the assessment of “helpfulness” into the following four progressive perspectives.
  1. 1.
    Validity measures whether the generated text is a valid feedback text.
  2. 2.
    Contextualization measures whether the generated text is a feedback text specific to the current input.
  3. 3.
    Constructiveness measures whether the feedback text provides critiques or suggestions w.r.t. a certain criterion.
  4. 4.
    Helpfulness measures whether the feedback text is helpful for improving the text writing w.r.t. a certain criterion.
By sequentially applying these perspectives, we can efficiently filter out feedback that fails to meet lower-level perspectives, ensuring that only the most relevant and constructive feedback reaches the final stage of evaluation. This method significantly reduces the noise in the feedback that is ultimately assessed for helpfulness. We designed different prompt templates for evaluating each perspective (see Appendix Tab. 13~Tab. 24) and validated our model-based evaluation approach by running a small-scale human evaluation conducted by two experienced NLP researchers using the same prompts as annotation guidelines. For each perspective and task, we randomly chose 100 feedback samples from all models, totaling 300 samples per perspective for annotation. The results indicate that our model-based evaluation method is highly accurate, with average accuracies of 96.48% for validity (95.33% inter-annotator agreement), 94.23% for contextualization (92% inter-annotator agreement), 95.22% for constructiveness (91% inter-annotator agreement), and 88.53% for helpfulness (89.67% inter-annotator agreement). [^8] When calculating the accuracy, we only include the samples where both annotators agreed on the label. More details are in Appendix A.6.
@@ 5 Experiments
In the case of the Crit, ICL, and CrICL strategies, the resulting feedback text is a collection of feedback text for each of the aspects.
@@
Human Annotation Details
Both processes of deriving criteria and creating demonstrations require meticulous manual refinement. We provide details of this process involving human input below. For this study, this has been carried out by two experienced NLP researchers. In terms of the human annotators’ role in verifying the collection of criteria, their main responsibility is to identify and remove duplicates without making additional edits. With a dataset comprising three tasks, 25 aspects, and 83 criteria, the average number of criteria per aspect is approximately 3.3. The time required to review each criterion for duplication is about 10 seconds, resulting in a total deduplication time of roughly 30 minutes. For the creation of demonstration inputs and outputs, our approach involved 332 demonstrations across three tasks. On average, each demonstration required about 8 minutes of curation and editing for both the input and the output. This effort amounted to a total of approximately 44 hours of human annotation.
@@ 6.1 Validity and Contextualization
Our study begins by examining LLMs’ capabilities in generating valid and contextual feedback texts for given examples. Specifically, we calculate the percentage of feedback texts that are valid and the percentage of valid texts that are contextual. The detailed results are in Appendix Tab. 25. Our key findings are as follows: (i) Regardless of the strategy used to generate the feedback, these LLMs almost always generate valid feedback, demonstrating their strong instruction-following and in-context learning capability. (ii) The inclusion of demonstrations comes with the risk of distracting models into writing feedback to the input text of a demonstration, sometimes resulting in a lower overall contextualization of the generated feedback. This is especially true if the demonstrations are long (e.g., introduction task).
@@ 6.2 Constructiveness and Helpfulness
For feedback texts that passed the validity and contextualization checks, we further scrutinize their constructiveness by counting the percentage of criteria in §4 they touched upon by providing critiques or suggestions. For feedback texts that passed the constructiveness check as well, we examine their helpfulness by counting the percentage of those critiques or suggestions that are helpful for improvement. Results are in Tab. 1.
@@
(i) For all three tasks across all models, the inclusion of criteria generally increases the constructiveness of generated feedback considerably (on average 25.8%, at most 44.6%) , outperforming the base strategy. (ii) In most cases, adding demonstrations improves constructiveness of generated feedback compared to the base strategy (on average 10.1%, at most 27.0%) . Only in a few cases does the addition of demonstrations lead to a noticeable decrease in constructiveness. This is observed in the introduction task, where the Command and LongAlpaca models produce less constructive feedback with demonstrations, likely struggling with extremely lengthy demonstrations. (iii) Interestingly, when both demonstrations and criteria are applied, the outcome is mostly comparable to adding demonstrations alone. This may be attributed to the extended length of demonstrations, which could potentially dilute the impact of the criteria.
@@
(i) The effect of adding criteria on helpfulness is model dependent. For GPT4, adding criteria always helps to generate more helpful critiques compared to the base strategy. For the other models, on the other hand, in all tasks, adding only criteria typically results in a marginally lower rate of helpful critiques, as compared to the base strategy. (ii) In most cases, adding demonstrations could improve the helpfulness compared to the base strategy (on average 2.7%, at most 13.7%). There are a few exceptions, such as when the base strategy already produces very helpful critiques (over 90% helpfulness). (iii) In most cases, combining both demonstrations and criteria yields inferior results compared to using demonstrations alone. This may be attributed to the criteria’s lack of specific guidance on crafting helpful critiques and its potential interference with the efficacy of demonstrations when added to the context.
@@ 6.3 Overall Performance
We also provide the overall performance of each model using different strategies. In particular, we count the total number of criteria the generated feedback texts have addressed by providing helpful critiques or suggestions for the 100 test samples for each task. The results are shown in Tab. 2. We find that (i) Providing criteria will always improve the overall performance regardless of the model and task. (ii) Providing demonstrations will generally improve the overall performance except for the cases where the overly long demonstrations negatively affect the contextualization of feedback too much (in the case of introduction task for most models). (iii) Adding both criteria and demonstrations typically does not outperform adding criteria alone, mostly because models generate fewer critiques/suggestions when demonstrations are provided compared to criteria only (see Appendix Tab. 26). This may be due to the fact that the feedback texts in demonstrations can include positive comments, thus causing the model-generated feedback to be less critical. Therefore, in practice, providing criteria only to LLMs is a reasonably good strategy for feedback generation.
@@
Table 2: Overall performance of each model using different strategies in terms of the number of criteria the generated feedback texts addressed through providing helpful critiques or suggestions.
@@ 7.1 Fine-grained Analysis of Criteria and Demonstration Effects
We selected the code task as our primary testbed, owing to its extensive number of criteria. For each criterion, we analyzed the ratio of generated critiques out of the total critiques that could be generated by the model for 100 test samples. We specifically excluded feedback with validity or contextualization errors. The percentage of generated critiques for each criterion is depicted on the left in Tab. 3. Furthermore, we examined the proportion of helpful critiques within the set of critiques for each criterion, with these percentages presented on the right in Tab. 3.
We observe that: (i) adding criteria enhances the model’s ability to generate critiques across various criteria. Notably, the effectiveness of this enhancement is more pronounced for criteria with which the model is already somewhat familiar, as evidenced by a higher baseline percentage of critiques. (ii) Adding demonstrations typically results in a more uniform distribution of helpful critiques across all criteria, indicating an overall improvement in critique quality. This approach contrasts with the base strategy, which tends to overlook certain criteria.
@@
To explore the effect of providing criteria at the granularity of an aspect and at the granularity of a single criterion, we chose the Reddit post writing task where each aspect has an average of 4.2 criteria. We only compare the feedback generated by the strategy of offering criteria only. The results are in Tab. 4. We see that providing criteria in an aspect’s granularity almost always positively affects the validity and contextualization of the generated feedback. However, providing criteria in an aspect’s granularity negatively affects the constructiveness of generated feedback, while the impact on helpfulness is model-dependent. In summary, providing criteria at the granularity of one aspect rather than on an individual criterion does not significantly affect overall performance, while offering greater efficiency during inference time.
@@ 8 Conclusion
In this paper, we present a general framework for teaching LLMs to use criteria for feedback generation that takes the guidelines as a starting point, semi-automatically extracts criteria from them, and constructs in-context demonstrations. We then apply these criteria and in-context demonstrations to guide the feedback generation process of LLMs. In order to provide a more comprehensive assessment of the quality of feedback, we propose a layered evaluation methodology to measure the quality of generated feedback from different perspectives. Experiments conducted on three writing tasks and seven different LLMs provide insights on the most effective way of teaching LLMs to use criteria. We also analyzed the effectiveness of applying criteria at different levels of granularity and found that providing criteria at the granularity of an aspect is more effective than providing criteria at the granularity of a single criterion.
@@ 9 Limitations and Future Work
In our experimental setting, we only consider LLMs that have strong instruction-following capabilities and a large context window. However, there are models that are not as strong at instruction following and in-context learning as the models we have chosen. How to teach those models to use criteria would be an interesting future work. In addition, our evaluation is primarily a model-based approach, which may introduce some inaccuracies, although the meta-evaluation shows good correlation with human judgments. Developing better model-based methods to measure feedback quality is also an important direction to explore in the future.
@@ A.5 Prompts For Layered Evaluation
A.6 Meta-Evaluation for Our Evaluation Methodology
While our model-based evaluation demonstrates high accuracy in alignment with human judgments, we have also undertaken a detailed analysis of human label distribution within the annotation task to gain deeper insights into these accuracy metrics. Our analysis reveals a notable imbalance in the distribution of "yes" labels, with high percentages observed for validity and contextualization (both at 93%), in contrast to a lower percentage for constructiveness (39%) and a moderate percentage for helpfulness (82%).
To ensure a more balanced and thorough assessment of the model’s evaluation capabilities across these dimensions, we have computed macro F1 scores. The results are as follows: validity at 83.38%, contextualization at 70.24%, constructiveness at 95.05%, and helpfulness at 77.25%.
A.7 Detailed Experiment Results
A.7.1 Results on Validity and Contextualization